OpenStudy (anonymous):

What are the dimensions of a rectangle that has a perimeter of 74 units and an area of 102 square units? a. 18 by 19 c. 6 by 17 b. 2 by 51 d. 3 by 34

OpenStudy (solomonzelman):

A - Area P - Perimeter l - length w - width \(\Huge\color{blue}{ \bf A= w \times l}\) \(\Huge\color{blue}{ \bf P= 2(w + l)}\) plug in your number \(\Huge\color{blue}{ \bf 102= w \times l}\) \(\Huge\color{blue}{ \bf 74= 2(w + l)}\) can you solve for w and l ? doesn't matter which one is which.

OpenStudy (solomonzelman):

you can just plug in your choices into the perimeter equation and into the Area equation see which of the choices works. Alternatively, you can solve the system using substitution. Are you able to do that?

OpenStudy (kc_kennylau):

Don't reply with "so", tell us whether you understand it or not, and ask us questions if you have any